OpenStudy (anonymous):

write with positive exponents: (x^-2y / z^-3)^3

hero (hero):

Hint: \[a^{-b} = \frac{1}{a^b}\]

OpenStudy (anonymous):

\[(\frac{ 1 }{ x^2y } / \frac{ 1 }{ z^3 })^3 \]

OpenStudy (anonymous):

i think :/

hero (hero):

You only inverse what is in the exponent. And when you inverse the fraction, whatever was in the exponent remains in the exponent. You don't bother inversing any variable that doesn't have a negative number for an exponent.

OpenStudy (anonymous):

im still confused with this stuff can you help me and walk me through how to do this one?